OpenStudy (anonymous):

Solve the perfect square trinomial equation. 9x^2+6x+1=4

OpenStudy (anonymous):

Please help!!!!!!!

OpenStudy (anonymous):

isnt this history?

OpenStudy (unklerhaukus):

minus four from both sides,

OpenStudy (unklerhaukus):

then use the quadratic formula

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

Factor:\[(3x+3)(3x-1)=0\]Therefore either 3x + 3 = 0 or 3x - 1 = 0. Solve both linear equations.